OpenStudy (anonymous):

Solve the following quadratic trig equations algebraically (without graphing technology) for the domain 0 ≤ x ≤ 360°. Use exact solutions whenever possible. 6cos2(θ) + cos(θ) - 1 = 0 2 - 2sin2(θ) = cos(θ)Solve the following quadratic trig equations algebraically (without graphing technology) for the domain 0 ≤ x ≤ 360°. Use exact solutions whenever possible. A. 6cos2(θ) + cos(θ) - 1 = 0 B. 2 - 2sin2(θ) = cos(θ)

OpenStudy (anonymous):

Found the solution, no worries. For those of you looking for it: A. 6cos(Theta) + cos(Theta) - 1 = 0 [2cos(Theta) + 1] [3cos(Theta) - 1] = 0 2cos(Theta) + 1 = 0 or 3cos(Theta) - 1 = 0 cos(Theta) = -1/2 or cos(Theta) = 1/3 (Theta )=180-60 or (Theta)=180+60 or (Theta)=70.5 or (Theta)=360-70.5 (Theta) = 120 or (Theta) = 240 or (Theta) = 70.5 or (Theta) = 289.5 b. 2 - 2sin(Theta) = cos(Theta) 2[sin(Theta) + cos(Theta)] - 2sin(Theta) = cos(Theta) 2sin(Theta) + 2cos(Theta) - 2sin(Theta) = cos(Theta) 2cos(Theta) - cos(Theta) = 0 cos(Theta) [2cos(Theta) - 1] = 0 cos(Theta) = 0 or cos(Theta) = (Theta) = 90 or (Theta) = 270 or (Theta) = 60 or (Theta) = 360-60 (Theta) = 90 or (Theta) = 270 or (Theta) = 60 or (Theta) = 300 --------------------------------------------------------- Found the solution here: http://hk.knowledge.yahoo.com/question/question?qid=7008071400443